OpenStudy (anonymous):

What is the value of tan(sin^-1(1/2))

OpenStudy (linn99123):

let's use the angle sum formulas for sin and cos, and the definition that tan = sin / cos tan (a + b) = sin (a + b) / cos (a + b) sin (a + b) = [sin a cos b + cos a sin b] cos (a + b) = [cos a cos b - sin a sin b] let a = sin^-1(-sqrt(3) / 2), so sin a = -sqrt(3) / 2 at this angle (in Q4), cos a = 1/2 (using 30-60-90 triangle, or the fact that sin^2 + cos^2 = 1) let b = cos^-1 (-sqrt(3) / 2), so cos b = -sqrt(3) / 2 at this angle (in Q2), sin b = 1/2 thus, tan (a + b) = [-sqrt(3)/ 2 (-sqrt(3)/ 2) + (1/2)(1/2)] / [-sqrt(3) / 2 (1/2) - (-sqrt(3) / 2)(1/2)] tan (a + b) = [3/4 + 1/4] / [-sqrt(3) / 4 + sqrt(3) / 4] tan (a + b) = 1 / (0), and is thus undefined tan (sin^-1(-sqrt(3) / 2) + cos^-1(-sqrt(3) / 2)) is undefined this is the same answer I got using the angle sum formula for tan(a + b) = (tan a + tan b) / [1 - tan a tan b] sin a = -sqrt(3) / 2 ==> tan a = -sqrt(3) cos b = -sqrt(3) / 2==> tan b = -sqrt(3) / 3 (tan a + tan b) / (1 - tan a tan b) denominator = 1 - tan a tan b = 1 - (-sqrt(3))(-sqrt(3) / 3) = 1 - 3/3 = 1 - 1 = 0 tan (a + b) is undefined let's verify with the angles (since we can) sin^-1 (-sqrt(3) / 2) = -60 cos^-1 (-sqrt(3) / 2) = 150 150 - 60 = 90, and yes, tan 90 is undefined

OpenStudy (zarkon):

let \(\theta=\sin^{-1}(1/2)\) then \(\sin(\theta)=1/2\) draw a right triangle 

OpenStudy (zarkon):

use Pythagorean theorem to get the 3rd side

OpenStudy (zarkon):



OpenStudy (zarkon):

\(\tan(\theta)\) is the opposite side over the adjacent side therefore \[\tan(\sin^{-1}(1/2))=\tan(\theta)=\frac{1}{\sqrt{3}}\]